# Microscopic Analysis on LLM players via Social Deduction Game

### v1

@@ III-C Data Collection
A total of 168 gameplay logs were collected. In this experiment, we adopted seven hidden locations from the original game. For each location, we ran 24 games: 2 types of citizens $\times$ 4 spy LLMs $\times$ 3 trials. Table I lists the statistics of the collected gameplay logs. On average, each game took 5.4 turns, with 23.8 reasoning steps. Here, the reasoning step means a unit of LLM’s reasoning for taking an action.
@@
IV Multifaceted Quantitative Analysis
Our study aimed to conduct a multifaceted quantitative analysis of LLM performance in SDGs. We designed seven metrics that enable fine-grained analysis of two key gameplay facets: intent identification and camouflage. Compared with previous studies that utilized single-faceted metrics, our approach allows for an analysis of intermediate gameplay behaviors. In this section, we discuss whether existing single-faceted metrics are sufficient for measuring intermediate behaviors. Next, we demonstrate a multifaceted quantitative analysis using the proposed metrics to explore gameplay behavior.
@@ IV-A Insufficiency of Previous Analysis
Previous studies commonly adopted single-faceted analyses to examine SDG gameplay behavior[17, 16, 2, 1]. The most frequently used metrics are the winning rates (WR) and living rounds (LR). These metrics focus on the final state of the game to measure the performance of a player: WR utilizes winner information, while LR uses the number of turns played.
  • Winning Rate (WR) is the percentage of games won by the spy out of the total number of games.
  • Living Round (LR) is the average number of turns played across all games.
@@ IV-A1 Metric result
Tables II (a) and III present the results of the single-faceted metrics when playing against strong and weak citizens, respectively. The WR and LR exhibited different trends in SDG play performance. In terms of WR, GPT-4 showed the highest performance against strong citizens (80.95%), followed by Gemini (57.14%), LLaMA2 (52.38%), and GPT-3.5 (33.33%). A similar trend was observed for weak citizens. However, in terms of LR, LLaMA2 demonstrated the highest performance against strong citizens (9.00), followed by Gemini (6.95), GPT-4 (5.80), and GPT-3.5 (3.71). This trend is similar when examining the performance against weak citizens. Thus, these metrics yield different results regarding the best-performing model: WR indicates GPT-4 as the best, while LR identifies LLaMA2 as the best.
This discrepancy is due to multiple winning conditions specified in the SDG. These conditions allow players to win games in several ways. Therefore, to accurately identify the skills that contribute to winning, it is necessary to consider multiple abilities simultaneously. However, since WR and LR consider only the final state of the game, these metrics cannot examine different abilities separately. For example, in SpyGame, the LR metric indicates how well a spy survives during the game. However, a longer survival duration does not always guarantee victory, leading to potential discrepancies between LR and WR results. Hence, a multifaceted method is essential for a comprehensive analysis of gameplay performance.
@@
(a) General (b) Intent identification (c) Camouflage
Winning Living Guess Information Notice Caught Vote Vote
Model Rate (%) Round (#) Success (%) Deduction (%) Rate (%) Rate (%) Rate (%) Entropy
Gemini 57.14 6.95 50.00 30.00 33.33 022.22 35.7 0.348
@@
(a) General (b) Intent identification (c) Camouflage
Winning Living Guess Info. Notice Caught Vote Vote
Model Rate (%) Round (#) Success (%) Catching (%) Rate (%) Rate (%) Rate (%) Entropy
Gemini 90.48 5.28 85.71 61.11 090.48 00.00 07.1 0.420
@@
IV-B Intent identification
We designed multiple metrics to enable a multifaceted analysis of intent identification and camouflage. We operationally define Intent identification as the ability to uncover hidden information through deduction or by capturing pieces of information from the conversation. Intent identification has two sub-abilities: “information capture” and “information deduction.” Information capture mainly focuses on how well a spy grabs crucial hints that a citizen might inadvertently reveal. Since strong citizens rarely make such mistakes, it is highly challenging to evaluate this sub-ability against them. On the other hand, information deduction (ID) mainly focuses on how effectively a spy infers a hidden location by combining various pieces of information. Since weak citizens usually provide information that directly exposes the hidden location, it is not meaningful to assess this sub-ability against weak citizens. Therefore, metrics should consider whether the spy is effectively utilizing the information provided by these citizens. To evaluate both information capture and deduction, we analyzed outcomes of the guessing by the spy. The following subsections detail the design of the four metrics and their results. The results indicated that GPT-4 was the most powerful spy player, followed by Gemini, GPT-3.5, and LlaMA2.
@@ IV-B1 Metric design
Based on a spy’s guesses during a game, we suggest four metrics for assessing intent identification. Four metrics range from relatively summative to highly formative.
  • Guess Success (GS) measures how much the guessing of a spy contributed to their victory. This metric focuses only on games where the spy attempted to guess, rather than all played games. We calculate GS by dividing the number of games in which the spy made a successful guess by the number of games in which guessing was attempted. A higher GS indicates that the spy has successfully identified hidden locations, demonstrating strong intent identification skills.
  • Notice Rate (NR) indicates how well a spy noticed the hidden location across all games. It is calculated by computing the percentage of games in which the spy correctly detected the hidden location over the total number of games played. Unlike GS, which is measured only when the spy decides to reveal their identity publicly, NR can be computed across all games. This is because NR mainly considers whether the spy “notices” the game regardless of whether they decide to guess or not.
  • Information Catching (IC) measures how often the spy successfully leverages exposed information to achieve a victory. Unlike NR, which only assesses the ability of the spy to identify the hidden location, IC evaluates the effectiveness in employing such information to win. We specifically focus on games where weak citizens directly exposed to information about the hidden location directly, instead of considering all played games. We measured IC by computing the proportion of games in which the spy successfully guessed the hidden location guessing from those where exposure occurred. Thus, a high IC score reflects the ability of the spy to achieve victory by effectively capturing and utilizing crucial information.
@@
In particular, we devised a multifaceted interpretation of intent identification by using both IC and ID metrics simultaneously. By combining the IC and ID with their average values, three combinations were derived. The first combination is the most balanced, where both IC and ID are high. This represents an ideal scenario where the spy successfully identifies the hidden location regardless of the type of citizen. This signifies that the model is good at both sub-abilities: information capture and information deduction. The second combination occurs when only one of the metrics, either IC or ID, is higher than the average. This suggests a lack of proficiency in the corresponding sub-ability. For instance, a model with high IC but low ID demonstrates strong information capture but weak information deduction. The final and least favorable combination for intent identification is when both IC and ID are low. This combination indicates that the spy is unable to infer the hidden location in most games.
Moreover, by combining GS and NR, we can explore whether the spy effectively incorporates the incoming information to maintain their guess. The ideal combination is when both GS and NR are high. This indicates that the spy consistently maintained a correct guess after identifying the hidden location, until the end of the game. In contrast, a high GS paired with a low NR suggests that while the spy struggled to predict the hidden location initially, he/she manages to maintain their guess once the correct location is identified. Meanwhile, the combination of low GS and high NR implies that although the spy successfully guessed the hidden location, the spy is prone to discarding the guess easily. The worst-case scenario is when both NR and GS are low, indicating a significant failure in both identifying and maintaining the correct hidden location.
@@ IV-B2 Result
Tables II (b) and III (b) present the measurements for intent identification when playing against strong and weak citizens, respectively. The results indicate that GPT-4 has sufficient intent identification ability compared to the other models. GPT-4 was the best model in terms of GS, followed by Gemini, GPT-3.5, and LlaMA2 (81.25, 50.0, 27.78, and 0%, respectively). The results in terms of NR show a similar trend: GPT-4 (61.90%), Gemini (33.33%), GPT-3.5 (28.57%), and LlaMA2 (9.52%). Finally, regarding ID, GPT-4 was the best model, followed by Gemini, GPT-3.5, and LlaMA2 (61.90, 30.0, 28.57, and 0%, respectively). In summary, GPT-4 demonstrated the best performance against strong citizens in intent identification. Meanwhile, LlaMA2 showed the lowest performance against strong citizens.
Against weak citizens, GPT-4 was the best model in terms of GS, followed by Gemini, GPT-3.5, and LlaMA2 (90.48, 85.71, 71.43, and 60%, respectively). The NR results displayed the same trend: GPT-4 (100%), Gemini (90.48%), GPT-3.5 (76.19%), and LlaMA2 (71.43%). Lastly, regarding IC, GPT-3.5 (92.86%) showed the best performance, differing in GS and NR, followed by GPT-4 (87.5%), Gemini (61.11%), and LlaMA2 (20.0%). Similar to the results against strong citizens, GPT-4 showcased good performance against weak citizens.
@@ IV-B3 Discussion
In this section, we discuss how to derive insights using the proposed metrics to conduct a multifaceted intent identification analysis. This discussion encompasses three aspects: (1) a complementary relationship between IC and ID; (2) an exploration of the spy’s filtering behavior on incoming information; and (3) the need for formative metrics concerning intent identification.
@@
First, ID and IC are complementary metrics that enable multifaceted analysis of intent identification. As aforementioned, these two metrics can be combined to form three distinct partitions, each representing different levels of intent identification ability. Using these partitions, we can determine the best model possessing sufficient sub-abilities for intent identification; In our analysis, GPT-4 demonstrated high scores in both ID and IC, with its ID being double that of the second-place model. Meanwhile, GPT-3.5 falls within the high-IC region but not the high-ID region. This indicates that while GPT-3.5 struggles to combine hints from the conversation to deduce the hidden location.
@@
Second, by analyzing GS and NR, we can examine how the spy filters incoming information to maintain their guess. As suggested in the metric design, we can use four partitions made by GS and NR. Among the four models, GPT-4 belongs to the first partition and achieves a high score on both metrics. This indicates that GPT-4 can filter useful information, reinforcing the belief of the spy in the hidden location. This could be a hint that GPT-4 has strong information deduction sub-ability. Meanwhile, belonging to different partitions can induce differences in such filtering behaviors. For example, although Gemini and GPT-3.5 exhibited similar NR scores against strong citizens, the GS of Gemini was almost double that of GPT-3.5. This difference suggests that Gemini is more selective in using guess actions than GPT-3.5, as evidenced by the fact that Gemini used 14 guess actions regardless of citizens, whereas GPT-3.5 utilized 21 (against weak) and 18 (against strong) guess actions.
@@
Finally, our intent identification metrics enable more detailed analyses of the factors contributing to the scores of simple metrics. Let us consider the simple metric, Win- ning Rate (WR), on Gemini models. Against strong citizens, Gemini showed a similar WR to LlaMA2. Meanwhile, against weak citizens, Gemini exhibited the same WR as GPT-4. However, these observations are somewhat shallow when we consider our intent identification metrics, as the three models display different intent identification abilities. For example, against strong citizens, Gemini had a higher ID score than LLaMA2 in that LLaMA2 completely failed to combine clues about the hidden location. Similarly, against weak citizens, we can conclude that GPT-4 is slightly better than Gemini because of its higher GS and NR scores. Thus, proposed metrics can reveal finer differences between LLMs than existing metrics.
@@
In addition, such a detailed analysis revealed that GPT-3.5 is overly confident in its deduction results. As illustrated in Tables II and III, the spy with GPT-3.5 demonstrated relatively lower NR scores; however, the spy used guessing actions more frequently. That is, the spy rushed into making guesses even when their inference about the hidden location was incorrect. One possible explanation is that GPT-3.5 may treat its own inference as hard evidence, as discussed in Section V. We suspect that this mechanism may have contributed to GPT-3.5 achieving the lowest WR score, despite its relatively good intent identification ability.
IV-C Camouflage
@@
In addition to intent identification, we designed four metrics for the multifaceted analysis of camouflage. Here, we operationally define camouflage as the ability to blend in with others during gameplay without revealing one’s true identity. Camouflage encompasses two sub-abilities: “evading suspicion” and “dispersing suspicion.” Evading suspicion mainly considers how well the spy avoids being suspected by others. It is quantified by the number of suspicions directed at the spy. On the other hand, dispersing suspicion primarily considers how difficult the spy makes it for all citizens to converge on a single suspect. It is quantified by the probability distribution of suspicions). To estimate the number and distribution of suspicions, we utilized logs of Day/Final votes. The following subsections describe the design of the four metrics and their results. The results indicate that GPT-3.5 is the most powerful spy player, followed by GPT-4, Gemini, and LlaMA2.
@@ IV-C1 Metric design
Based on day/final votes and accusations during a game, we suggest four metrics for measuring camouflage, ranging from relatively summative to highly formative.
  • Caught Rate (CR) is how frequently the spy is caught by citizens in the Day/Final vote. We compute this by dividing the number of games in which the spy was caught by the number of games lost. Thus, a high CR indicates that the majority of citizens easily identified the spy player during the game. Note that this is a summative metric for the spy’s camouflage performance because it only considers the final state of the game.
  • Accusation Rate (AR) measures how frequently citizens accuse the spy. We measure this by computing the proportion of correct accusations made across all games. Unlike CR, which summarizes all citizens’ thoughts, AR focuses on whether an individual citizen can strongly suspect a true spy since the turn leader is the only player who can make an accusation. Depending on the gameplay, it is possible that there is no accusation during the game; in this case, AR cannot be measured.
@@
We devised a multifaceted interpretation of camouflage by using VR and VE simultaneously. By partitioning VR and VE by their average values, we derived four combinations. First, the most balanced combination was a low VR and high VE. This is an ideal combination because the spy can successfully evade others’ suspicions and lead citizens to discourage each other from narrowing down spy candidates. Second, models in which VR and VE are both low or high lack one of the two sub-abilities. Finally, the worst camouflage combination occurs when VR is high and VE is low. This combination indicates that citizens’ suspicions were concentrated on the spy and that the spy failed to undermine suspicions.

### v2

@@ III-B Models and Data
To collect data, we let each tested model play SpyGame as a spy against strong or weak citizens. Here, the strong or weak citizen models were fixed based on performance reported in prior studies [21, 22, 1]: GPT-4 for strong and Llama2-70b-chat for weak citizens. And, we conducted our experiment repeatedly. Specifically, we used seven locations borrowing from the original game. For each location, we ran 24 games4442 types of citizens $\times$ 4 spy LLMs $\times$ 3 trials.. Consequently, we collected a total of 168 gameplay logs. Table I lists statistics of collected logs. On average, each game took 5.4 turns, with 23.8 reasoning steps which refers to a unit of LLM’s reasoning for taking an action.
@@
IV Microscopic Quantitative Analysis
Our study aimed to investigate whether LLMs can support obscured communication by conducting a microscopic analysis of their SDG playing. We designed seven metrics that enable fine-grained analysis of two key SDG facets: subtext inference and deceptive control. Compared with most metrics adopted by previous research, which were designed on game-level outcomes, our approach allows for an analysis of intermediate behaviors during the game. Also, our quantitative metrics are fully automated metrics that do not require any manual labeling. In this section, we first discuss whether existing simple metrics are sufficient for measuring intermediate behaviors of LLMs. Then, we demonstrate our analysis can discover new insights that could not be captured by using current metrics by discussing our quantitative results.
@@ IV-A Insufficiency of Previous Analysis
Previous studies commonly adopted macroscopic analyses to examine SDG gameplay behavior[8, 20, 19, 7, 6]. The most frequently used metrics are the win rates (WR) and living rounds (LR). These metrics evaluate a player’s performance based on final game outcomes: WR is determined by the win/loss result, while LR reflects how long the player remained active during the game.
  • Win Rate (WR) is the percentage of games won by the spy out of the total number of games.
  • Living Round (LR) is the average number of turns the spy survived, across all games.
@@ IV-A1 Metric result
Tables II (a) and III (a) present the results of the coarse-grained metrics when playing against strong and weak citizens, respectively. The WR and LR exhibited different trends in SDG play performance. In terms of WR, GPT-4 showed the highest performance against strong citizens (80.95%), followed by Gemini (57.14%), Llama (52.38%), and GPT-3.5 (33.33%). A similar order of performance was observed for against weak citizens. However, in terms of LR, Llama demonstrated the highest performance against strong citizens (9.00), followed by Gemini (6.95), GPT-4 (5.80), and GPT-3.5 (3.71). This trend is similar when examining the performance playing against weak citizens. Note that these metrics yield different results regarding the best-performing model: WR indicates GPT-4 as the best, while LR identifies Llama.
This discrepancy is due to multiple winning conditions specified in the SDG. These conditions allow players to win games in several ways. Therefore, to accurately identify the skills that contribute to winning, it is necessary to consider multiple abilities simultaneously. However, since WR and LR consider only the final state of the game, these metrics cannot examine different abilities separately. For example, in SpyGame, LR indicates how well a spy survives during the game. However, a longer survival does not always guarantee victory, leading to potential discrepancies between LR and WR. Hence, a microscopic method is required for a comprehensive analysis of LLM’s performance in various facets.
@@
(a) General (b) Subtext inference (c) Deceptive Control
Winning Living Guess Notice Information Information Caught Vote Vote
Model Rate (%) Round (#) Success (%) Rate (%) Catching Deduction (%) Rate (%) Rate (%) Entropy
Gemini 57.14 6.95 50.00 $3$ 30.00 022.22 35.7 0.348
@@
Note that Information Catching was not measured playing against strong citizens.
@@
(a) General (b) Subtext inference (c) Deceptive control
Winning Living Guess Notice Information Information Caught Vote Vote
Model Rate (%) Round (#) Success (%) Rate (%) Catching (%) Deduction Rate (%) Rate (%) Entropy
Gemini 90.48 5.28 85.71 090.48 61.11 - 00.00 07.1 0.420
@@
Note that Information Deduction was not measured playing against weak citizens.
IV-B Subtext Inference
We designed multiple metrics to enable a microscopic analysis of subtext inference. We operationally define subtext inference as the ability to uncover hidden information through deduction or by capturing pieces of information from the conversation. Subtext inference has two sub-abilities: information capture and information deduction. Information capture measures whether a spy notices and exploits critical hints inadvertently revealed by citizens. Since weak citizens generally leak such information unintentionally, information capture ability becomes essential for the spy to win the game against them. Since strong citizens rarely make such mistakes, it is highly challenging to observe information capture of a spy. On the other hand, information deduction mainly focuses on how effectively a spy infers a hidden location by combining various pieces of information. Since strong citizens generally conceal the key information, information deduction ability becomes essential for the spy to uncover information embodied in a citizen’s utterance.
To evaluate both information capture and deduction, we analyzed the outcomes of the spy’s guessing. The following subsections detail the design of the four metrics and their results. The results indicated that GPT-4 was the most powerful spy player, followed by Gemini, GPT-3.5, and Llama.
@@ IV-B1 Metric design
To assess the spy’s ability to identify hidden intent, we propose four metrics based on their guessing behavior. These metrics range from relatively summative to more formative.
  • Guess Success (GS) measures how much the guessing of a spy contributed to their victory. This metric focuses only on games where the spy attempted to guess, rather than all played games. We calculate GS by dividing the number of games in which the spy made a successful guess by the number of games in which guessing was attempted. A higher GS indicates that the spy has successfully identified hidden locations, demonstrating strong subtext inference skills.
  • Notice Rate (NR) indicates how well a spy noticed the hidden location across all games. It is calculated by computing the percentage of games in which the spy correctly detected the hidden location over the total number of games played. Unlike GS, which is measured only when the spy decides to reveal their identity publicly, NR can be computed across all games. This is because NR mainly considers whether the spy notices the hidden location, regardless of whether they decide to guess.
  • Information Catching (IC) measures how often the spy successfully leverages exposed information to achieve a victory. Unlike NR, which only assesses the ability of the spy to identify the hidden location, IC evaluates the effectiveness in employing such information to win. We specifically focus on games where weak citizens directly exposed information about the hidden location instead of considering all the played games. Here, the exposure means a case where a citizen’s utterance literally contains the hidden location [^5] Qualitative analysis in Section V-B1 discusses the case.. We measured IC by computing the proportion of games in which the spy successfully guessed the hidden location from games where exposure occurred. Thus, a high IC score reflects the ability of the spy to achieve victory by effectively capturing and utilizing key information.
@@
We provide a multifaceted interpretation of subtext inference by jointly analyzing IC and ID metrics. Based on whether each metric is above or below its average, we classify model behavior into three combinations. The first combination, when both IC and ID are high, represents an ideal scenario in which the spy successfully identifies the hidden location regardless of the type of citizen. The second combination occurs when only one of the metrics, either IC or ID, is higher than the average. This suggests a relative lack of proficiency in the corresponding sub-ability. For instance, a model with high IC but low ID demonstrates strong information capture but weak information deduction. The final and least favorable combination for subtext inference is when both IC and ID are low. This combination indicates that the spy failed to infer the hidden location in most games.
Moreover, by combining GS and NR, we can explore whether the spy effectively incorporates the incoming information to maintain their guess. The ideal combination is when both GS and NR are high. This indicates that the spy consistently maintained a correct guess after identifying the hidden location, until the end of the game. In contrast, a high GS paired with a low NR suggests that while the spy struggled to predict the hidden location initially, they managed to maintain their guess once the correct location was identified. Meanwhile, the combination of low GS and high NR implies that although the spy successfully guessed the hidden location, the spy is prone to discarding the guess easily. The worst-case scenario is when both NR and GS are low, indicating a significant failure in both identifying and maintaining the correct hidden location.
@@ IV-B2 Result
Tables II (b) and III (b) present the measurements for subtext inference when playing against strong and weak citizens, respectively. The results indicate that GPT-4 has sufficient ability compared to the other models. GPT-4 was the best model in terms of GS, followed by Gemini, GPT-3.5, and Llama (81.25, 50.0, 27.78, and 0%, respectively). The results in terms of NR show a similar trend: GPT-4 (61.90%), Gemini (33.33%), GPT-3.5 (28.57%), and Llama (9.52%). Finally, regarding ID, GPT-4 was the best model, followed by Gemini, GPT-3.5, and Llama (61.90, 30.0, 28.57, and 0%). In summary, GPT-4 demonstrated the best performance against strong citizens in subtext inference. Meanwhile, Llama showed the lowest performance against strong citizens.
Against weak citizens, GPT-4 was the best model in terms of GS, followed by Gemini, GPT-3.5, and Llama (90.48, 85.71, 71.43, and 60%, respectively). NR displayed the same trend: GPT-4 (100%), Gemini (90.48%), GPT-3.5 (76.19%), and Llama (71.43%). Lastly, regarding IC, GPT-3.5 (92.86%) showed the best performance, differing in GS and NR, followed by GPT-4 (87.5%), Gemini (61.11%), and Llama (20.0%). Similar to the results against strong citizens, GPT-4 showcased good performance against weak citizens.
@@ IV-B3 Discussion
In this section, we derive insights using the proposed metrics by conducting a multifaceted analysis. This discussion encompasses three aspects: (1) a complementary relationship between IC and ID; (2) an exploration of the spy’s filtering behavior on incoming information; and (3) the need for formative metrics concerning subtext inference.
@@
First, IC and ID are complementary metrics that enable multifaceted analysis of subtext inference. As aforementioned, these two metrics can be combined to form three distinct partitions, each representing different levels of subtext inference ability. Using these partitions, we can determine the best model possessing sufficient sub-abilities for subtext inference; In our analysis, GPT-4 demonstrated high scores in both IC and ID, with its ID being double that of the second-place model. Meanwhile, GPT-3.5 falls within the high-IC region but not the high-ID region. This indicates that while GPT-3.5 struggles to combine hints from the conversation to deduce the hidden location.
@@
Second, by analyzing GS and NR, we can examine how the spy filters incoming information to maintain their guess. As suggested in the metric design, we can use four quadrants made by GS and NR. Among the four models, GPT-4 belongs to the first partition and achieves a high score on both metrics. This indicates that GPT-4 can filter useful information, reinforcing the belief of the spy in the hidden location. This could be a hint that GPT-4 has strong information deduction sub-ability. Meanwhile, belonging to different partitions can induce differences in such filtering behaviors. For example, although Gemini and GPT-3.5 exhibited similar NR scores against strong citizens, GS of Gemini was almost double that of GPT-3.5. This difference suggests that Gemini is more selective in using guess actions than GPT-3.5, as evidenced by the fact that Gemini used 14 guess actions regardless of citizens, whereas GPT-3.5 utilized 21 (against weak) and 18 (against strong) guess actions.
@@
Finally, our metrics regarding subtext inference enable more detailed analyses of the factors contributing to the scores of simple metrics. Let us consider the simple metric, Winning Rate (WR), on Gemini models. Against strong citizens, Gemini showed a similar WR to Llama. Meanwhile, against weak citizens, Gemini exhibited the same WR as GPT-4. However, these observations are somewhat shallow when we consider our metrics, as the three models display different subtext inference abilities. For example, against strong citizens, Gemini had a higher ID score than Llama in that Llama completely failed to combine clues about the hidden location. Similarly, against weak citizens, we can conclude that GPT-4 is slightly better than Gemini because of its higher GS and NR scores. Thus, proposed metrics can reveal finer differences between LLMs than existing metrics.
@@
In addition, such a detailed analysis revealed that GPT-3.5 is overly confident in its deduction results. As illustrated in Tables II and III, the spy with GPT-3.5 demonstrated relatively lower NR scores; however, the spy used guessing actions more frequently. That is, the spy rushed into making guesses even when their inference about the hidden location was incorrect. One possible explanation is that GPT-3.5 may treat its own inference as hard evidence, as we will discuss in Section V. We suspect that this mechanism may have contributed to GPT-3.5 achieving the lowest WR score, despite its relatively good subtext inference ability.
IV-C Deceptive Control
@@
In addition to subtext inference, we designed four metrics for the multifaceted analysis of deceptive control. Here, we operationally define deceptive control as the ability to blend in with others by controlling information. Deceptive control encompasses two sub-abilities: evading suspicion and dispersing suspicion. Evading suspicion mainly considers how well the spy avoids being suspected by others. It is quantified by the number of suspicions directed at the spy. On the other hand, dispersing suspicion primarily considers how difficult the spy makes it for all citizens to converge on a single suspect. It is quantified by the probability distribution of suspicions. To estimate the number and distribution of suspicions, we utilized logs of Day/Final votes. The following subsections describe the design of the four metrics and their results. The results indicate that GPT-3.5 is the most powerful spy player, followed by GPT-4, Gemini, and Llama.
@@ IV-C1 Metric design
Based on day/final votes and accusations during a game, we suggest three metrics for measuring deceptive control, ranging from relatively summative to highly formative.
  • Caught Rate (CR) is how frequently the spy is caught by citizens in the Day/Final vote. We compute this by dividing the number of games in which the spy was caught by the number of games lost. Thus, a high CR indicates that the majority of citizens easily identified the spy player during the game. Note that this is a summative metric for the spy’s deceptive control ability because it only considers the final state of the game.
@@
We devised a multifaceted interpretation of deceptive control by using VR and VE simultaneously. By partitioning VR and VE by their average values, we derived four combinations. First, the most balanced combination was a low VR and high VE. This is an ideal combination because the spy can successfully evade others’ suspicions and lead citizens to discourage each other from narrowing down spy candidates. Second, models in which VR and VE are both low or high lack one of the two sub-abilities. Finally, the worst combination occurs when VR is high and VE is low. This combination indicates that citizens’ suspicions were concentrated on the spy and that the spy failed to undermine suspicions.